OpenStudy (anonymous):

Write an equation of a line with a slope of 2/3 and y-intercept of 7. Write the equation in both slope intercept form and standard form. 7y=2/3x+? ?????

OpenStudy (anonymous):

y-intercept is when the line is on the y-axis, right? so that's when x = 0. So we have this : y = mx+b 7 = 2/3*0 + b 7 = 0 +b y = 2/3x + 7

OpenStudy (anonymous):

your y-intercept is quite conveniently always b